# From Complex-Analytic Models to Sparse Domination: A Dyadic Approach of Hypersingular Operators via Bourgain’s Interpolation Method

### v4

@@ Abstract.
Motivated by the work of Cheng–Fang–Wang–Yu on the hypersingular Bergman projection, we develop a real-variable and dyadic framework for hypersingular operators in regimes where strong-type estimates fail at the critical line. The main new input is a hypersingular sparse domination principle combined with Bourgain’s interpolation method, which provides a flexible mechanism to establish critical-line (and endpoint) estimates.
@@
including a restricted weak-type bound at $(p,q)=\bigl(\tfrac{1}{3-2t},1\bigr)$. Finally, we introduce a class of hypersingular cousin of sparse operators in $\mathbb{R}^{n}$ associated with graded sparse families, quantified by the sparseness $\eta$ and a new structural parameter (the degree) $K_{\mathcal{S}}$, and we characterize the corresponding strong/weak/restricted weak-type regimes in terms of $(n,t,\eta,K_{\mathcal{S}})$.
Our real-variable perspective addresses an inquiry raised by Cheng–Fang–Wang–Yu on developing effective real-analytic tools in the hypersingular regime for $K_{2t}$, and it also provides a new route toward the critical-line analysis of Forelli–Rudin type operators and related hypersingular operators in both real and complex settings.
@@
Hypersingular maximal operator, Hypersingular Bergman projection, Hypersingular sparse operator, graded family, critical line estimates
@@ 3. $\bullet$
A common feature of these results is that they provide strong $L^{p}$$L^{q}$ bounds only away from the critical line, while estimates on the critical line itself appear to be unavailable in the literature. One main reason is that strong-type bounds on the critical line generally fail (see, Figure 2). This obstruction, in turn, make it difficult to use the techniques that are effective in the case $t\leq 1$, since $K_{2t}$ becomes more singular in the hypersingular regime (see the discussion after (1.4)). Moreover, for general hypersingular operators, weak-type bounds may even fail at one of the endpoints on the critical line (see Remark 1.8), indicating that genuinely new ideas are required to treat critical-line estimates.
@@
In [14, Section 1], the authors remarked that, in the hypersingular regime $t>1$, developing a workable real-variable approach that yields explicit boundedness results for concrete operators remains an open direction. This is one of the main motivations for the present work.
The first goal of the current paper is to develop harmonic-analytic methods for the study of the hypersingular Bergman projections $K_{2t}$ for $t>1$, thanks to the recent development in dyadic harmonic analysis. In particular, we aim to establish estimates on the critical line, which, to the best of our knowledge, are even new in the unit disc setting. In this way, we address an inquiry raised in [14] by developing a real–variable approach that yields explicit bounds in the hypersingular regime.
@@
In what follows, we will restrict our attention to the dyadic hypersingular maximal operator $\mathcal{M}_{t}^{\mathcal{D}}$. It is a standard fact (via the $1/3$-trick) that $\mathcal{M}_{t}$ is pointwise comparable to the sum of two dyadic counterparts: there exist two dyadic systems $\mathcal{D}$ and $\widetilde{\mathcal{D}}$ on $\mathbb{T}$ such that
@@
Here $\mathcal{D}$ may be taken to be the standard dyadic system on $\mathbb{T}$, and $\widetilde{\mathcal{D}}$ a $1/3$-shifted dyadic system. This reduction goes back to Mei’s work on $BMO$ and its dyadic analog [27].
@@
    Condition (1.3) can also be interpreted from the perspective of Bergman–Carleson embeddings in complex function theory, initiated in a series of influential works of Luecking [22, 23, 24]. In that setting, the boundedness of the embedding operator $id:A^{p}({\mathbb{D}})\to L^{q}_{\mu}({\mathbb{D}})$ (here, $A^{p}({\mathbb{D}})$ denotes the standard Bergman space) is characterized by an $L^{\frac{p}{p-q}}({\mathbb{D}})$-integrability condition on an appropriate testing function. From this viewpoint, the above result also suggests a way to understand the Békollé–Bonami condition through Carleson embedding.
@@
    (see, [14, Theorem 3] and also Theorem 5.7, (1) for a generalization and strengthening of this result in terms of “hypersingular” sparse operators);
@@
[…]
    (see, Lemma 4.1 and Corollary 4.5).
  3. (3)
    (Restricted weak-type bound) bounded from $L^{p,1}({\mathbb{D}})$ to $L^{q,\infty}({\mathbb{D}})$ for
    (1.4) $p=\frac{1}{3-2t}\quad\textrm{and}\quad q=1.$
    (see, Theorem 4.4).
@@
Figure 2. Boundedness of $K_{2t}$ for $1<t<3/2$: the red line and the shaded region indicate strong $(p,q)$ bounds, the blue line indicates weak $(p,q)$ bounds, and the green dot indicates restricted $(p,q)$ bounds.
One major difficulty in analyzing $K_{2t}$ arises at the endpoint $p=\frac{1}{3-2t}$ and $q=1$. Indeed, as observed in [14, Theorem 3], strong-type estimates fail on the critical line 1/q1/p=2t21/q-1/p=2t-2. This is in sharp contrast with the Bergman projection $K_{2}$, whose $L^{2}$ boundedness is simply guaranteed by its definition. As a consequence, both the Calderón–Zygmund approach (for instance, the Calderón–Zygmund decomposition, which uses an $L^{r}$ bound, $1<r<\infty$, to control the bad function) and more recent sparse domination techniques (e.g. the weak-type machinery in [9, Theorem E]) become more delicate to implement directly in this setting.
To address this issue, we develop a new approach based on Bourgain’s interpolation trick (see, Lemma 4.3).
Key idea. Observe that the point $\left(1/p,1/q\right)=(3-2t,1)$ lies between the two corner points $(0,1)$ and $(1,1)$, which correspond to a “good” strong-type regime and a “bad” blow-up regime, respectively. Our strategy is to connect these two behaviors by combining a dyadic/sparse domination perspective with Bourgain’s interpolation trick. Remarkably, this interpolation lands exactly at $\left(1/p,1/q\right)=(3-2t,1)$ and, together with off-diagonal Marcinkiewicz interpolation, yields weak-type bounds along the critical line away from this endpoint.
@@
    (1.5) $K_{2t}^{+}f(z):=\int_{{\mathbb{D}}}\frac{f(w)}{|1-z\overline{w}|^{2t}}\,dA(w),$
@@
  2. (2)
    One may ask whether, at the endpoint (1/p,1/q)=(32t,1)(1/p,1/q)=(3-2t,1), the restricted weak-type estimate (1.4) can be upgraded to a weak-type bound (recall that $\mathcal{M}_{t}^{\mathcal{D}}$ does satisfy a weak-type estimate at this endpoint). The situation, however, is more subtle. Indeed, one can show that the hypersingular Berezin transform (1.5) is unbounded from $L^{\frac{1}{3-2t}}({\mathbb{D}})$ to $L^{1,\infty}({\mathbb{D}})$ (see, Proposition 4.2). For $K_{2t}$, we conjecture that the weak-type estimate also fails at this endpoint (see Section 6 for further discussion).
Motivated by our approach the critical-line estimate for $K_{2t}$, it is natural to ask whether Bourgain’s interpolation method can be used to study hypersingular analogues of sparse operators in harmonic analysis. This leads to the second main goal of the present paper. More precisely, let $t>1$ and let $\mathcal{S}$} be a sparse family in $\mathbb{R}^{n}$ such that [^2] This global containment assumption is natural in the hypersingular setting (see Remark 5.3 for further discussion). there exists a dyadic cube $Q_{0}$ with $Q\subseteq Q_{0}$ for all $Q\in{\mathcal{S}}$. Consider the following hypersingular averaging operator
@@
We are interested in determining the admissible $(p,q)$-range for which $\mathbb{A}_{{\mathcal{S}}}^{t}$ is of strong type, weak type, or restricted weak type. Note that the study of $\mathbb{A}_{{\mathcal{S}}}^{t}$ is also of independent interest from the viewpoint of dyadic harmonic analysis, since it can be viewed as a hypersingular counterpart of the classical sparse operator.
@@
Here is the main result in the second part of the paper. Let $\mathcal{S}$} be a sparse family in $\mathbb{R}^{n}$ as above, with sparseness $\eta\in(0,1)$ and degree $K_{{\mathcal{S}}}\in[1,\infty)$; we refer to such a family as a graded sparse family. Then for any $1<t<1-\frac{\log_{2}(1-\eta)}{nK_{{\mathcal{S}}}}$, we have $\mathbb{A}_{{\mathcal{S}}}^{t}$ is
@@
[…]
  3. (3)
    (Restricted weak-type bound) bounded from $L^{p,1}({\mathbb{D}})$ to $L^{q,\infty}({\mathbb{D}})$ for
    $p=\frac{-\log_{2}(1-\eta)}{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}\quad\textrm{and}\quad q=1.$
    However, in general the endpoint weak-type estimate for $\mathbb{A}_{{\mathcal{S}}}^{t}$ fails. Indeed, this follows from Proposition 4.2 together with the fact that $K_{2t}^{+}$ can be bounded above pointwisely by a sum of two hypersingular sparse operators (see (4.1)).
@@
Figure 3. Boundedness of $\mathbb{A}_{{\mathcal{S}}}^{t}$ for $1<t<1-\frac{\log_{2}(1-\eta)}{nK_{{\mathcal{S}}}}$: the red line and the shaded region indicate strong $(p,q)$ bounds, the blue line indicates weak $(p,q)$ bounds, and the green dot indicates restricted $(p,q)$ bounds.
@@
    A noteworthy feature of the main theorems above is that the geometric parameters of the family—in particular, the sparseness $\eta$ and the degree $K_{{\mathcal{S}}}$—enter the admissible $(p,q)$-range through the critical relation
@@
    The results above strengthen our earlier characterization of the hypersingular Bergman projection $K_{2t}$ in the range $1<t<3/2$. Indeed, the collection of all dyadic Carleson boxes on $\mathbb{D}$} (in real dimension $n=2$) forms a graded sparse family with sparseness $\eta=1/2$ and degree $K_{{\mathcal{S}}}=1$. With these parameters, one has
@@
    The strong-type range above is sharp, as can be seen from the strong-type characterization for $K_{2t}$ (see [14, Theorem 3]). This suggests that the $L^{p}$$L^{q}$ mapping properties of $K_{2t}$ are governed primarily by the underlying geometry of the (dyadic) Carleson boxes, rather than by analyticity.
@@
    From the viewpoint of complex function theory, we develop a systematic theory of the hypersingular maximal operator ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ and also advance the study of Forelli–Rudin type operators in the hypersingular regime by establishing new endpoint and critical-line estimates for $K_{2t}$ with $t>1$. To the best of our knowledge, these estimates are new even in the unit disc setting. A key feature of our approach is the introduction of Bourgain’s interpolation method into this setting, which allows us to overcome the lack of strong-type bounds and provides a robust harmonic-analytic framework for hypersingular operators arising in complex variables.
@@
    More broadly, the Bourgain interpolation framework developed in this paper provides a flexible and systematic tool for treating hypersingular operators. We expect that this approach will have further applications, including weighted estimates, commutator theory, hypersingular Bergman theory in several complex variables (on general domains), and related problems, both in complex analysis and in other areas of analysis; see Section 6 for possible directions for future work.
The rest of this paper is organized as follows. In Section 2, we collect basic definitions and notation that will be used throughout the paper. In Section 3, we establish the $L^{p}$ mapping theory for the dyadic hypersingular maximal operator $M_{t}^{\mathcal{D}}$, including both the off-critical-line bounds and the critical-line estimates, together with the endpoint characterizations in the radial weighted setting. Moreover, we prove a two-weight estimate for ${\mathcal{M}}_{t}^{{\mathcal{D}}},t>0$. In Section 4, we turn to the hypersingular Bergman projection $K_{2t}$ and prove the critical-line estimates, with special emphasis on the endpoint $(p,q)=\left(\frac{1}{3-2t},1\right)$ via a combination of hypersingular sparse domination and Bourgain’s interpolation trick. In Section 5, we introduce hypersingular sparse operators associated with graded sparse families in $\mathbb{R}^{n}$ and prove the corresponding $L^{p}$-$L^{q}$ theory, which in particular extends and strengthens the results obtained earlier for $K_{2t}$. Finally, in Section 6, we propose several open problems and further directions motivated by the methods developed in this paper.

### v5

# From Complex–Analytic Models to Dyadic Methods: A Real–Variable Approach to Hypersingular Operators

@@ Abstract.
Motivated by the work of Cheng–Fang–Wang–Yu on the hypersingular Bergman projection, we develop a real-variable, dyadic framework for hypersingular operators in regimes where strong-type bounds fail on the critical line. Our main new input is what we call the Forelli–Rudin method: a dyadic real-variable mechanism, inspired by complex-analytic Forelli–Rudin type arguments, that provides a flexible way to obtain critical-line and endpoint estimates for general hypersingular operators.
@@
with particular emphasis on the weak-type bound at the endpoint $(p,q)=\bigl(\tfrac{1}{3-2t},1\bigr)$. Finally, we introduce a class of hypersingular cousin of sparse operators in $\mathbb{R}^{n}$ associated with graded sparse families, quantified by the sparseness $\eta$ and a new structural parameter (the degree) $K_{\mathcal{S}}$, and we characterize the corresponding strong and weak type regimes in terms of $(n,t,\eta,K_{\mathcal{S}})$.
Our real-variable perspective addresses an inquiry raised by Cheng–Fang–Wang–Yu on developing effective real-analytic tools in the hypersingular regime for both ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ and $K_{2t}$, and it also provides a new route toward the critical-line analysis of Forelli–Rudin type operators and related hypersingular operators in both real and complex settings.
@@
Hypersingular maximal operator, Hypersingular Bergman projection, Hypersingular sparse operator, graded family, critical line estimates, Forelli–Rudin method
@@ 3. $\bullet$
A common feature of these results is that they provide strong $L^{p}$$L^{q}$ bounds only away from the critical line, while estimates on the critical line itself appear to be unavailable in the literature. One main reason is that strong-type bounds on the critical line generally fail (see, Figure 2). This obstruction, in turn, make it difficult to use the techniques that are effective in the case $t\leq 1$, since $K_{2t}$ becomes more singular in the hypersingular regime.
@@
In [14, Section 1], the authors remarked that in the hypersingular regime $t>1$, it remains an open direction to develop a real-variable approach that yields explicit boundedness results for concrete operators. This is one main motivation for the present work.
The first goal of the current paper is to develop harmonic-analytic methods for the study of the hypersingular Bergman projections $K_{2t}$ for $t>1$, thanks to the recent development in dyadic harmonic analysis. In particular, we aim to establish estimates on the critical line, which, to the best of our knowledge, are even new in the unit disc setting.
@@
In what follows, we will restrict our attention to the dyadic hypersingular maximal operator $\mathcal{M}_{t}^{\mathcal{D}}$. It is a standard fact (via the $1/3$-trick) that $\mathcal{M}_{t}$ is pointwise comparable to the sum of two dyadic counterparts: there exists two dyadic systems $\mathcal{D}$ and $\widetilde{\mathcal{D}}$ on $\mathbb{T}$ such that [^2] Here we refer to $\mathcal{D}$} and $\widetilde{{\mathcal{D}}}$ as a pair of adjacent dyadic systems on $\mathbb{T}$.
@@
Here $\mathcal{D}$ may be taken to be the standard dyadic system on $\mathbb{T}$, and $\widetilde{\mathcal{D}}$ is the $1/3$-shifted dyadic system. This reduction dates back to the work of Garnett, Jones, and Mei on $BMO$ and its dyadic counterpart (see [16, 28]).
@@
    Condition (1.3) can also be interpreted from the perspective of Bergman–Carleson embeddings in complex function theory, initiated in a series of influential works of Luecking [23, 24, 25]. In that setting, the boundedness of the embedding operator $id:A^{p}({\mathbb{D}})\to L^{q}_{\mu}({\mathbb{D}})$ (here, $A^{p}({\mathbb{D}})$ denotes the standard Bergman space on $\mathbb{D}$}) is characterized by an $L^{\frac{p}{p-q}}({\mathbb{D}})$-integrability condition on an appropriate testing function. From this viewpoint, the above result also suggests a way to understand the Békollé–Bonami condition through Carleson embedding.
@@
    (see, [14, Theorem 3], and also Theorem 5.7, (1) for a generalization and strengthening of this result in terms of hypersingular sparse operators);
@@
    $\frac{1}{q}-\frac{1}{p}=2t-2,\quad\textrm{with}\ 1\leq p,q\leq+\infty.$
    (see, Lemma 4.1, Proposition 4.2, and Corollary 4.4).
@@
Figure 2. Boundedness of $K_{2t}$ for $1<t<3/2$: the red line and the shaded region indicate strong $(p,q)$ bounds, and the blue line indicates weak $(p,q)$ bounds.
One main difficulty in analyzing $K_{2t}$ occurs at the endpoint $p=\frac{1}{3-2t}$ and $q=1$ (see, Proposition 4.2). As noted in [14, Theorem 3], strong-type bounds fail along the critical line 1/q1/p=2t21/q-1/p=2t-2, so $K_{2t}$ behaves more singularly in this regime. This is in sharp contrast with the Bergman projection $K_{2}$, whose $L^{2}$ boundedness is is simply guaranteed by its definition.
To address this issue, we introduce a new approach, which we call the Forelli–Rudin method. For the endpoint behavior of $K_{2t}$ at $(p,q)=\bigl(\tfrac{1}{3-2t},1\bigr)$, we will use its complex–analytic version, which serves as a motivating case.
Key idea 1: A complex–analytic version of the Forelli–Rudin method (motivating case). To treat $K_{2t}$ at the endpoint $(p,q)=\bigl(\tfrac{1}{3-2t},1\bigr)$, the main idea is to use the factorization
(1.4) $K_{2t}f(z)=W(z)\,B_{2t}f(z),\qquad z\in{\mathbb{D}},$
where $W(z)=(1-|z|^{2})^{2-2t}$ and $B_{2t}f(z):=(1-|z|^{2})^{2t-2}K_{2t}f(z)$. This separates the singularity into a simple weight $W$ and a less singular Forelli–Rudin type operator $B_{2t}$, thereby reducing the endpoint weak-type bound to a strong-type estimate for $B_{2t}$ together with a borderline Lorentz control of $W$.
@@
$K_{2t}^{+}f(z):=\int_{{\mathbb{D}}}\frac{f(w)}{|1-z\overline{w}|^{2t}}\,dA(w),$
@@
Motivated by our approaches to ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ and $K_{2t}$, it is natural to ask whether these methods can be extended to study hypersingular analogues of sparse operators in harmonic analysis. This leads to the second main goal of the present paper. More precisely, let $t>1$ and let $\mathcal{S}$} be a sparse family in $\mathbb{R}^{n}$ such that [^3] This global containment assumption is natural in the hypersingular setting (see Remark 5.3 for further discussion). there exists a dyadic cube $Q_{0}$ with $Q\subseteq Q_{0}$ for all $Q\in{\mathcal{S}}$. Consider the following hypersingular averaging operator
@@
We are interested in determining the admissible $(p,q)$-range for which $\mathbb{A}_{{\mathcal{S}}}^{t}$ is of strong type, or weak type. Note that the study of $\mathbb{A}_{{\mathcal{S}}}^{t}$ is also of independent interest from the viewpoint of dyadic harmonic analysis, since it can be viewed as a hypersingular counterpart of the classical sparse operator.
@@
Here are the main results in the second part of the paper. Let $\mathcal{S}$} be a sparse family in $\mathbb{R}^{n}$ with sparseness $\eta\in(0,1)$ and degree $K_{{\mathcal{S}}}\in[1,\infty)$; we refer to such a family as a graded sparse family. Then for any $1<t<1-\frac{\log_{2}(1-\eta)}{nK_{{\mathcal{S}}}}$, we have $\mathbb{A}_{{\mathcal{S}}}^{t}$ is
@@
    $\frac{1}{q}-\frac{1}{p}=\frac{nK_{{\mathcal{S}}}(t-1)}{-\log_{2}(1-\eta)},\quad\textrm{with}\ 1\leq p,q\leq+\infty.$
@@
Figure 3. Boundedness of $\mathbb{A}_{{\mathcal{S}}}^{t}$ for $1<t<1-\frac{\log_{2}(1-\eta)}{nK_{{\mathcal{S}}}}$: the red line and the shaded region indicate strong $(p,q)$ bounds, and the blue line indicates weak $(p,q)$ bounds.
The crux of the analysis for $\mathbb{A}_{{\mathcal{S}}}^{t}$ is to establish the weak-type estimate at the endpoint
(1.5) $p=\frac{-\log_{2}(1-\eta)}{-\log_{2}(1-\eta)+nK_{{\mathcal{S}}}(1-t)}\qquad\text{and}\qquad q=1.$
The difficulties at this endpoint are twofold:
  • As for $K_{2t}$, strong-type bounds fail along the critical line for $\mathbb{A}_{{\mathcal{S}}}^{t}$. Consequently, both the Calderón–Zygmund approach (for instance, the Calderón–Zygmund decomposition, which relies on an $L^{r}$ bound, $1<r<\infty$, to control the bad function) and more recent sparse domination techniques (e.g. the weak-type machinery in [9, Theorem E]) are more delicate to implement in this hypersingular setting.
    Indeed, we present two independent proofs, one based on sparse domination machinery and the other using Bourgain’s interpolation trick, showing that at the endpoint (1.5) the operator $\mathbb{A}_{{\mathcal{S}}}^{t}$ satisfies a restricted weak-type estimate. While this is weaker than the weak-type bound proved in Theorem 5.7, it already suffices to obtain weak-type estimates on the critical line away from this endpoint. We refer to Section 5.3 for details.
  • In this dyadic setting, we also lack an underlying complex-analytic geometry. More precisely, in several complex variables, quantities such as 1|z|1-|z| are typically interpreted as the boundary distance  $\hbox{ \rm dist}(z,\mathbb{T})$. In contrast, for $\mathbb{A}_{{\mathcal{S}}}^{t}$ there is no canonical notion of boundary distance, and it is unclear what should play the role of the “boundary” of the region.
Key idea 2: A dyadic version of the Forelli–Rudin method. To exploit the idea in (5.14), the key observation is that the boundary distance 1|z|1-|z| admits a convenient dyadic re-encoding. Fix $z\in{\mathbb{D}}$ and let $\mathcal{D}$} be any dyadic system on $\mathbb{T}$. Define
${\bf N}(z):=\#\{\,I\in{\mathcal{D}}:\ z\in Q_{I}\,\},\qquad z\in{\mathbb{D}},$
namely, the number of dyadic Carleson boxes $Q_{I}$ (with $I\in{\mathcal{D}}$) that contain $z$. Then
${\bf N}(z)\simeq\log_{2}\frac{1}{1-|z|},$
and hence one may view
$1-|z|\simeq 2^{-{\bf N}(z)}.$
Using now this idea for $\mathbb{A}_{{\mathcal{S}}}^{t}$, one can prove the pointwise bound
(1.6) $\mathbb{A}_{{\mathcal{S}}}^{t}f(x)\lesssim{\bf W}(x)\,M^{{\mathcal{D}}}_{\mathrm{HL}}f(x),\qquad x\in\mathbb{R}^{n},$
where $\bf W$} is a suitable weight function and $M^{{\mathcal{D}}}_{\mathrm{HL}}$ denotes the dyadic Hardy–Littlewood maximal operator on $\mathbb{R}^{n}$ associated to $\mathcal{D}$}. (1.6) is precisely the dyadic analogue of (1.4) (see Theorem 5.7, Step III for more details). The weak-type bound at the endpoint (1.5) then follows by establishing the appropriate Lorentz-space estimates for $\bf W$}.
We make some further remarks.
@@
    A noteworthy feature of Theorem 5.7 is that the geometric parameters of the family—in particular, the sparseness $\eta$ and the degree $K_{{\mathcal{S}}}$—enter the admissible $(p,q)$-range through the critical relation
@@
    The results above strengthen our earlier characterization of the hypersingular Bergman projection $K_{2t}$ in the range $1<t<3/2$. Indeed, this follows from the pointwise sparse bounds
    $\left|K_{2t}f(z)\right|\lesssim\mathbb{A}_{\{Q_{I}\}_{I\in{\mathcal{D}}}}^{t}f(z)+\mathbb{A}_{\{Q_{I}\}_{I\in\widetilde{{\mathcal{D}}}}}^{t}f(z),\qquad z\in{\mathbb{D}},$
[…]
@@
    The strong-type range above is sharp, as can be seen from the strong-type characterization for $K_{2t}$ (see [14, Theorem 3]). These observations suggest that the $L^{p}$$L^{q}$ mapping properties of $K_{2t}$ are governed primarily by the geometry of (dyadic) Carleson boxes, rather than analyticity.
@@
    From the viewpoint of complex function theory, we develop a systematic theory of the hypersingular maximal operator ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ and also advance the study of Forelli–Rudin type operators in the hypersingular regime by establishing new endpoint and critical-line estimates for $K_{2t}$ with $t>1$. To the best of our knowledge, these estimates are new even in the unit disc setting. A key feature of our approach is the introduction of Forelli–Rudin method into this setting, which allows us to overcome the lack of strong-type bounds and provides a robust harmonic-analytic framework for hypersingular operators in both real and complex settings.
@@
    More broadly, the Forelli–Rudin methed developed in this paper provides a flexible and systematic tool for treating hypersingular operators. We expect that this approach will have further applications, including weighted estimates, commutator theory, hypersingular Bergman theory in several complex variables (on general domains), and related problems, both in complex analysis and in other areas of analysis; see Section 6 for possible directions for future work.
The rest of this paper is organized as follows. In Section 2, we collect basic definitions and notation that will be used throughout the paper. In Section 3, we establish the $L^{p}$ mapping theory for the dyadic hypersingular maximal operator $M_{t}^{\mathcal{D}}$, including both the off-critical-line bounds and the critical-line estimates, together with the endpoint characterizations in the radial weighted setting. Moreover, we prove a two-weight estimate for ${\mathcal{M}}_{t}^{{\mathcal{D}}},t>0$. In Section 4, we turn to the hypersingular Bergman projection $K_{2t}$ and prove the critical-line estimates, with special emphasis on the endpoint $(p,q)=\left(\frac{1}{3-2t},1\right)$ via a complex–analytic version of the Forelli–Rudin method. In Section 5, we introduce hypersingular sparse operators associated with graded sparse families in $\mathbb{R}^{n}$ and prove the corresponding $L^{p}$-$L^{q}$ theory via a dyadic version of the Forelli–Rudin method, which in particular extends and strengthens the results obtained earlier for $K_{2t}$. In addition, we include a discussion of two other different approaches to study $\mathbb{A}_{{\mathcal{S}}}^{t}$: one via ideas from sparse domination, and the other via Bourgain’s interpolation trick. Finally, in Section 6, we propose several open problems and further directions motivated by the methods developed in this paper.